OpenStudy (anonymous):

On a grid, the distance between the points located at (-3, -4) and (-8, 1) is roughly equal to _____.

OpenStudy (nurali):

Use the Pythagorean theorem. (-8-(-3))^2+(1-(-4))^2=distance^2 25+25=d^2 50=d^2 square root both sides d=7.07